# ToMCCA-3: A realistic 3-body coalescence model

### v1

@@ 4.3 Model uncertainties
In this section, the global yield uncertainties will be discussed. From the $A=2$ case [36], the uncertainty on the source size, caused by the measured deuteron and proton spectra uncertainty, was 5.7%. It was obtained by varying the $B$ parameter in Eq. 42 within the uncertainties obtained from fitting the deuteron spectra and is thus independent of $\langle m_{\mathrm{T}}\rangle$. When varying the source size by this amount, the 3He yield changed by ±$\pm 13\%$ for $\langle\mathrm{d}N_{\mathrm{ch}}/\mathrm{d}\eta\rangle_{|\eta|<0.5}$ $=2$ and by ±$\pm 18\%$ for $\langle\mathrm{d}N_{\mathrm{ch}}/\mathrm{d}\eta\rangle_{|\eta|<0.5}$ $=45$, giving on average a variation of ±$\pm 15.5\%$. Additionally, the cutoff momentum $\Lambda_{c}=1$ GeV/$c$ for the 3He and 3H wave functions was varied by ±$\pm 50$ MeV, which resulted in a global variation of ±$\pm 7\%$. Adding these two, a global uncertainty of ±$\pm 17\%$ is reached. In the case of ${}^{3}_{\Lambda}$, the cutoff variation can be omitted, and a global uncertainty of ±$\pm 15.5\%$ is taken.
@@ 5 Summary and Outlook
In this paper, we have presented an upgrade of the ToMCCA model to incorporate coalescence for $A=3$ nuclei, including 3He, 3H, and ${}^{3}_{\Lambda}$. The results for 3He and 3H are obtained for various different wave functions based on Argonne $\text{Argonne }v_{18}$ for the 2-body potential and Urbana IX for the genuine 3-body potential, as well as the Minnesota potential, which is a 2-body potential capable of reproducing the binding energies without additional 3-body potentials. The results for all wave functions that reproduce the binding energies are compatible with each other, and only the Argonne $\text{Argonne }v_{18}$ without UIX wave function shows a significant deviation. Interestingly, the 3H/3He ratio shows no dependence on the charged-particle multiplicity, which is in contrast to previous studies using Gaussian wave functions [58]. Overall, the description of 3He yields and their $p_{\mathrm{T}}$ spectra is in agreement with the measurements by ALICE in pp collisions at $\sqrt{s}=13$ TeV. In the case of ${}^{3}_{\Lambda}$ the Congleton approach [41] for the wave function was tested. It includes an undisturbed deuteron, described using the Argonne $\text{Argonne }v_{18}$ wave function, and a quasi-free $\Lambda$ around it. With this wave function, predictions for the yields and $p_{\mathrm{T}}$ spectra in pp collisions are presented. As of now, no data on ${}^{3}_{\Lambda}$ production in pp collisions has been published. The ${}^{3}_{\Lambda}$/3He ratio is predicted by ToMCCA to drop off with increasing $p_{\mathrm{T}}$, which is a common differentiator between thermal production and coalescence models. The minimum bias predictions of the ${}^{3}_{\Lambda}$ yield are about 3 times higher than the yield at the same mean multiplicity, but with a narrow multiplicity interval using Erlang functions.
With this successful upgrade of the ToMCCA model to $A=3$ coalescence it can be used for predictions of nuclei yields at different collision energies, specifically the ones interesting for indirect dark matter searches. Furthermore, the predictions could help in determining the production mechanism of light nuclei once high-precision data from the LHCs Run-3 and Run-4 campaigns become available.
Acknowledgments
@@
Appendix A 4D to 3D reduction
@@
In the so-called sudden approximation, the Lorentz-invariant yield of a trinucleon system of total spin s is given by:  [^6] Note that in these calculations, we use bold and italic fonts for three- and four-vectors. For example, the four-vector $x\equiv(t,{\bm{r}})$ and $p\equiv(p^{0},{\bm{p}})$. Moreover, the four-vectors product will be indicate with $p\cdot r\equiv p^{0}t-{\bm{p}}\cdot{\bm{r}}$
@@
$\gamma\frac{\mathrm{d}N_{\mathrm{A}}}{\mathrm{\leavevmode\nobreak\ d}^{3}P}=\frac{S_{\mathrm{A}}}{(2\pi)^{4}}\int\mathrm{\leavevmode\nobreak\ d}^{4}x_{1}\int\mathrm{\leavevmode\nobreak\ d}^{4}x_{2},\int\mathrm{\leavevmode\nobreak\ d}^{4}x_{3}\int\mathrm{\leavevmode\nobreak\ d}^{4}x_{1}^{\prime}\int\mathrm{d}^{4}x_{2}^{\prime}\int\mathrm{d}^{4}x_{3}^{\prime}\\ \times\Psi^{*}\left(x_{1}^{\prime},x_{2}^{\prime},x_{3}^{\prime},\right)\Psi\left(x_{1},x_{2},x_{3}\right)\rho_{1,2,3}\left(x_{1},x_{2},x_{3};x_{1}^{\prime},x_{2}^{\prime},x_{3}^{\prime}\right),$ (49)
@@
where $\Psi\left(x_{1},x_{2},x_{3}\right)$ is the three-particle bound state Bethe Salpeter amplitude and $\rho_{1,2,3}$ is the three-particle reduced density matrix, while the factor $S_{\mathrm{A}}$ accounts for the spin and isospin statistics of the considered case. The density matrix for three particle systems is assumed to be factored into single particle densities [Bellini:2020cbj] $\rho_{1,2,3}\left(x_{1},x_{2},x_{3};x_{1}^{\prime},x_{2}^{\prime},x_{3}^{\prime}\right)=\rho_{1}\left(x_{1};x_{1}^{\prime}\right)\times\rho_{1}\left(x_{2};x_{2}^{\prime}\right)\times\rho_{1}\left(x_{3};x_{3}^{\prime}\right)$, and the single-particle density is written in terms of single particle Wigner function, $f_{1}^{W}$,
@@
$\gamma\frac{\mathrm{d}N_{\mathrm{A}}}{\mathrm{\leavevmode\nobreak\ d}^{3}P}=\frac{S_{\mathrm{A}}}{(2\pi)^{16}}\int\mathrm{d}^{4}p_{1}\int\mathrm{d}^{4}p_{2}\int\mathrm{d}^{4}p_{3}\int\mathrm{d}^{4}x_{1}\int\mathrm{d}^{4}x_{2}\,\int\mathrm{d}^{4}x_{3}\int\mathrm{\leavevmode\nobreak\ d}^{4}x_{1}^{\prime}\int\mathrm{d}^{4}x_{2}^{\prime}\int\mathrm{d}^{4}x_{3}^{\prime}\\ \times\Psi^{*}\left(x_{1}^{\prime},x_{2}^{\prime},x_{3}^{\prime},\right)\Psi\left(x_{1},x_{2},x_{3}\right)\,e^{ip_{1}\cdot\left(x_{1}-x_{1}^{\prime}\right)+ip_{2}\cdot\left(x_{2}-x_{2}^{\prime}\right)+ip_{3}\cdot\left(x_{3}-x_{3}^{\prime}\right)}\,\\ \times W_{\mathrm{A}}\left(p_{1},p_{2},p_{3},\frac{x_{1}+x_{1}^{\prime}}{2},\frac{x_{2}+x_{2}^{\prime}}{2},\frac{x_{3}+x_{3}^{\prime}}{2}\right)\,.$ (53)
Now, we introduce the center-of-mass four vector $R$ and a pair of Jacobi four vectors $\xi_{1}$, $\xi_{2}$, related to $x_{1},x_{2},x_{3}$ by some linear relation which depends on the masses of the three particles. In general, we have $x_{i}=R+w_{1,i}\xi_{1}+w_{2,i}\xi_{2}$. Moreover, $d^{4}x_{1}d^{4}x_{2}d^{4}x_{3}=|J^{-1}|^{4}d^{4}Rd^{4}\xi_{1}d^{4}\xi_{2}$. The conjugate four momenta of $R,\xi_{1},\xi_{2}$ are the total four momentum $P=p_{1}+p_{2}+p_{3}$ and a pair of Jacobi four momenta $q_{1},q_{2}$. We have $p_{1}\cdot x_{1}+p_{2}\cdot x_{2}+p_{3}\cdot x_{3}=P\cdot R+\xi_{1}\cdot q_{1}+\xi_{2}\cdot q_{2}$. In general, we find $p_{i}=c_{i}P+$ some linear combinations of $k_{1}$ and $k_{2}$, where $c_{1}+c_{2}+c_{3}=1$. Finally, $d^{4}p_{1}d^{4}p_{2}d^{4}p_{3}=|J|^{4}d^{4}Pd^{4}q_{1}d^{4}q_{2}$.
In general, the Bethe-Salpeter amplitude $\Psi_{A}$ can be factorized into the plane wave for center-of-mass motion with total momentum $P_{\mathrm{A}}$ and an internal wave function that depends only on two relative coordinates $\xi_{1}$, $\xi_{2}$ describing the motion of internal nucleons/$\Lambda$ relative to the center of mass.
@@
$\gamma\frac{\mathrm{d}N_{\mathrm{A}}}{\mathrm{d}^{3}P}=\frac{S_{\mathrm{A}}}{(2\pi)^{20}}|J^{-1}|^{4}\int\mathrm{d}^{4}P\int\mathrm{d}^{4}q_{1}\int\mathrm{d}^{4}q_{2}\int\mathrm{\leavevmode\nobreak\ d}^{4}R\int\mathrm{d}^{4}\xi_{1}\,\int\mathrm{d}^{4}\xi_{2}\int\mathrm{d}^{4}R^{\prime}\int\mathrm{d}^{4}\xi_{1}^{\prime}\int\mathrm{d}^{4}\xi_{2}^{\prime}\\ \times e^{-iP_{\mathrm{A}}\cdot(R-R^{\prime})}\varphi_{\mathrm{A}}^{*}\left(\xi_{1}^{\prime},\xi_{2}^{\prime}\right)\varphi_{\mathrm{A}}\left(\xi_{1},\xi_{2}\right)\,e^{iP\cdot\left(R-R^{\prime}\right)+iq_{1}\cdot\left(\xi_{1}-\xi_{1}^{\prime}\right)+iq_{2}\cdot\left(\xi_{2}-\xi_{2}^{\prime}\right)}\,\\ \times W_{\mathrm{A}}\left(p_{1},p_{2},p_{3},\frac{x_{1}+x_{1}^{\prime}}{2},\frac{x_{2}+x_{2}^{\prime}}{2},\frac{x_{3}+x_{3}^{\prime}}{2}\right)\,.$ (55)
@@
$\gamma\frac{\mathrm{d}N_{\mathrm{A}}}{\mathrm{\leavevmode\nobreak\ d}^{3}P}=\frac{S_{\mathrm{A}}}{(2\pi)^{16}}|J^{-1}|^{4}\int\mathrm{d}^{4}q_{1}\int\mathrm{d}^{4}q_{2}\int\mathrm{\leavevmode\nobreak\ d}^{4}S\int\mathrm{d}^{4}\xi_{1}\,\int\mathrm{d}^{4}\xi_{2}\int\mathrm{d}^{4}\xi_{1}^{\prime}\int\mathrm{d}^{4}\xi_{2}^{\prime}\\ \times\varphi_{\mathrm{A}}^{*}\left(\xi_{1}^{\prime},\xi_{2}^{\prime}\right)\varphi_{\mathrm{A}}\left(\xi_{1},\xi_{2}\right)\,e^{iq_{1}\cdot\left(\xi_{1}-\xi_{1}^{\prime}\right)+iq_{2}\cdot\left(\xi_{2}-\xi_{2}^{\prime}\right)}\,\\ \times W_{\mathrm{A}}\left(p_{1},p_{2},p_{3},\frac{x_{1}+x_{1}^{\prime}}{2},\frac{x_{2}+x_{2}^{\prime}}{2},\frac{x_{3}+x_{3}^{\prime}}{2}\right)\Bigr{|}_{P=P_{\mathrm{A}}}\,.$ (56)
For the moment everything is exact. We will now approximate this relation in order to obtain the three-dimensinal relation given in Eq. (6). In the following, we introduce the notation $S=(t_{S},\bm{S})$, $R=(t_{R},{\bm{R}})$, $\xi_{i}=(\tau_{i},\bm{\xi}_{i})$, $i=1,2$, and $x_{i}=(t_{i},{\bm{x}}_{i})$, $i=1,\ldots,3$, and similarly for the primed quantities. Clearly $t_{S}=(t_{R}+t_{R}^{\prime})/2$ and $t_{i}=t_{R}+$ linear combination of $\tau_{1}$ and $\tau_{2}$. For more details see Ref. [Bellini:2020cbj]
@@
    $\int{dq_{1}^{0}\over 2\pi}{dq_{2}^{0}\over 2\pi}e^{iq_{1}^{0}(\tau_{1}-\tau_{1}^{\prime})+iq_{2}^{0}(\tau_{2}-\tau_{2}^{\prime})}=\delta(\tau_{1}-\tau_{1}^{\prime})\delta(\tau_{2}-\tau_{2}^{\prime})$ (57)

### v2

@@ 4.3 Model uncertainties
In this section, the global yield uncertainties will be discussed. From the $A=2$ case [42], the uncertainty on the source size, caused by the measured deuteron and proton spectra uncertainty, was 5.7%. It was obtained by varying the $B$ parameter in Eq. (42) within the uncertainties obtained from fitting the deuteron spectra and is thus independent of $\langle m_{\mathrm{T}}\rangle$. When varying the source size by this amount, the 3He yield changed by $\pm 13\%$ for $\langle\mathrm{d}N_{\mathrm{ch}}/\mathrm{d}\eta\rangle_{|\eta|<0.5}$ =2=2 and by $\pm 18\%$ for $\langle\mathrm{d}N_{\mathrm{ch}}/\mathrm{d}\eta\rangle_{|\eta|<0.5}$ =45=45, giving on average a variation of $\pm 15.5\%$. In order to ensure numeric stability, the relative momenta of the considered nucleons cannot exceed $k_{c}=1$ GeV/$c$. In order to gauge the effect of this cutoff, its value for the 3He and 3H AV18+UIX wave functions was varied by $\pm 50$ MeV, which resulted in a global variation of $\pm 7\%$. Adding these two, a global uncertainty of $\pm 17\%$ is reached. In the case of $^{3}_{\Lambda}$H, the cutoff variation can be omitted since the solution is numerically much more stable, and a global uncertainty of $\pm 15.5\%$ is taken.
@@ 5 Summary and Outlook
[…]
With this successful upgrade of the ToMCCA model to $A=3$ coalescence, it can be used for predictions of nuclei yields at different collision energies, specifically the ones interesting for indirect dark matter searches. Furthermore, the predictions could help in determining the production mechanism of light nuclei once high-precision data from the LHCs Run-3 and Run-4 campaigns become available.
Acknowledgments We gratefully acknowledge the support of the INFN-Pisa computing center.
@@
Appendix A Deuteron Spectra with new source parameterization
Figure 5: (left) Comparison of ToMCCA with ALICE deuteron ($(d+\overline{d})/2$)-spectra using the new parameterization of the source and the EPOS phase space. (right) Comparison of $B_{2}$ predictions by ToMCCA with the measurements of ALICE for pp collisions at $\sqrt{s}=5$ TeV. The colored bands indicate the previously determined 4.6% uncertainty on the predicted deuteron spectra, stemming from the experimental uncertainty of the proton spectra.
Since the source is retuned using the EPOS qrq-r-phase space as input, its performance with deuterons needs to be reassessed. The main change comes from the fact that particles with small relative momenta also have relatively smaller distances, thus boosting coalescence probability. Overall, the deuteron spectra are well described by the newly tuned model within the experimental and model uncertainties, as shown in Fig. 5.
Appendix B $\Lambda$ spectra and parameterization
Fig. 6 (left) depicts the $\Lambda$ spectra measured by ALICE [2] in pp collisions at $\sqrt{s}=13$ TeV in 10 multiplicity classes between $\langle\mathrm{d}N_{\mathrm{ch}}/\mathrm{d}\eta\rangle_{|\eta|<0.5}$ =2.5 and $\langle\mathrm{d}N_{\mathrm{ch}}/\mathrm{d}\eta\rangle_{|\eta|<0.5}$ =25.75. The ToMCCA predictions are shown alongside the Levy-Tsallis fit to the experimental data. The parameters of this Levy-Tsallis fit are shown in the right panel for the various multiplicity classes. Overall a good agreement with the measurements is achieved.
[…]
Appendix C $B_{3}$ parameter
Fig. 7 shows the measured $B_{3}$ parameter defined as
$B_{A}(p_{\mathrm{T,p}})=\frac{1}{2\pi p_{\mathrm{T,A}}}\frac{\mathrm{d}^{2}N_{A}}{\mathrm{d}y\mathrm{d}p_{\mathrm{T,A}}}\bigg/\left(\frac{1}{2\pi p_{\mathrm{T,p}}}\frac{\mathrm{d}^{2}N_{p}}{\mathrm{d}y\mathrm{d}p_{\mathrm{T,p}}}\right)^{A}$ (49)
with $A=3$ and $p_{\mathrm{T,p}}=p_{\mathrm{T,A}}/A$ for $p_{\mathrm{T}}/A=0.73$ GeV/$c$. In the inset, the minimum bias collisions [^9] INEL¿0 by ALICE definitions. are enhanced for visual clarity. Overall, ToMCCA describes the measurements well, with a slight tension for minimum bias collisions which are, however covered within 2 standard deviations.
[…]
Appendix D Minimum bias multiplicity
Throughout this work, predictions for minimum bias collisions are made. This requires a description of the measured minimum bias distributions. For this, published distributions by the ALICE collaboration [13] for energies from $\sqrt{s}=2.76$ TeV to $\sqrt{s}=13$ TeV are fitted using a two-Erlang distribution Ansatz
$P(N_{\rm ch})=\alpha\frac{\lambda_{1}^{\kappa_{1}}N_{\rm ch}^{\kappa_{1}-1}\exp(-\lambda_{1}N_{\rm ch})}{\Gamma(\kappa_{1})}+(1-\alpha)\frac{\lambda_{2}^{\kappa_{2}}N_{\rm ch}^{\kappa_{2}-1}\exp(-\lambda_{2}N_{\rm ch})}{\Gamma(\kappa_{2})}\ .$ (50)
Figure 8 shows the obtained fit parameters as a function of the collision energy alongside parameterizations using power laws ($ax^{b}$). Figure 9 shows the measured multiplicity distributions alongside the parameterized ones.
Figure 8: Parameters used in the two Erlang functions used to describe the minimum bias multiplicity distributions as a function of collision energy.
Figure 9: Published multiplicity distributions by the ALICE collaboration for $\sqrt{s}=2.76,5.02,7,8$ and $13$ TeV alongside the output of the Two-Erlang parameterization.
Note that the measurements are done in a range $0.8<\eta<0.8$ and the transformation to $\langle\mathrm{d}N_{\mathrm{ch}}/\mathrm{d}\eta\rangle_{|\eta|<0.5}$ has to be made.
Appendix E 4D to 3D reduction
@@ Appendix D Minimum bias multiplicity
In the so-called sudden approximation, the Lorentz-invariant yield of a trinucleon system is given by  [^10] Note that in these calculations, we use bold and italic fonts for three- and four-vectors. For example, the four-vector $x\equiv(t,{\bm{r}})$ and $p\equiv(p^{0},{\bm{p}})$. Moreover, the four-vectors product will be indicate with $p\cdot r\equiv p^{0}t-{\bm{p}}\cdot{\bm{r}}$.:
@@
$\gamma\frac{\mathrm{d}N_{\mathrm{A}}}{\mathrm{\penalty 10000\ d}^{3}P}=\frac{S_{\mathrm{A}}}{(2\pi)^{4}}\int\mathrm{\penalty 10000\ d}^{4}x_{1}\int\mathrm{\penalty 10000\ d}^{4}x_{2},\int\mathrm{\penalty 10000\ d}^{4}x_{3}\int\mathrm{\penalty 10000\ d}^{4}x_{1}^{\prime}\int\mathrm{d}^{4}x_{2}^{\prime}\int\mathrm{d}^{4}x_{3}^{\prime}\\ \times\Psi^{*}\left(x_{1}^{\prime},x_{2}^{\prime},x_{3}^{\prime},\right)\Psi\left(x_{1},x_{2},x_{3}\right)\rho_{1,2,3}\left(x_{1},x_{2},x_{3};x_{1}^{\prime},x_{2}^{\prime},x_{3}^{\prime}\right),$ (51)
@@
where $\Psi\left(x_{1},x_{2},x_{3}\right)$ is the three-particle bound state Bethe Salpeter amplitude and $\rho_{1,2,3}$ is the three-particle reduced density matrix, while the factor $S_{\mathrm{A}}$ accounts for the spin and isospin statistics of the considered case. The density matrix for three-particle systems is assumed to be factored into single-particle densities $\rho_{1,2,3}\left(x_{1},x_{2},x_{3};x_{1}^{\prime},x_{2}^{\prime},x_{3}^{\prime}\right)=\rho_{1}\left(x_{1};x_{1}^{\prime}\right)\times\rho_{1}\left(x_{2};x_{2}^{\prime}\right)\times\rho_{1}\left(x_{3};x_{3}^{\prime}\right)$, and the single-particle density is written in terms of single particle Wigner function, $f_{1}^{W}$,
@@
$\gamma\frac{\mathrm{d}N_{\mathrm{A}}}{\mathrm{\penalty 10000\ d}^{3}P}=\frac{S_{\mathrm{A}}}{(2\pi)^{16}}\int\mathrm{d}^{4}p_{1}\int\mathrm{d}^{4}p_{2}\int\mathrm{d}^{4}p_{3}\int\mathrm{d}^{4}x_{1}\int\mathrm{d}^{4}x_{2}\,\int\mathrm{d}^{4}x_{3}\int\mathrm{\penalty 10000\ d}^{4}x_{1}^{\prime}\int\mathrm{d}^{4}x_{2}^{\prime}\int\mathrm{d}^{4}x_{3}^{\prime}\\ \times\Psi^{*}\left(x_{1}^{\prime},x_{2}^{\prime},x_{3}^{\prime},\right)\Psi\left(x_{1},x_{2},x_{3}\right)\,e^{ip_{1}\cdot\left(x_{1}-x_{1}^{\prime}\right)+ip_{2}\cdot\left(x_{2}-x_{2}^{\prime}\right)+ip_{3}\cdot\left(x_{3}-x_{3}^{\prime}\right)}\,\\ \times W_{\mathrm{A}}\left(p_{1},p_{2},p_{3},\frac{x_{1}+x_{1}^{\prime}}{2},\frac{x_{2}+x_{2}^{\prime}}{2},\frac{x_{3}+x_{3}^{\prime}}{2}\right)\,.$ (55)
Now, we introduce the center-of-mass four vector $R$ and a pair of Jacobi four vectors $\xi_{1}$, $\xi_{2}$, related to $x_{1},x_{2},x_{3}$ by some linear relation which depends on the masses of the three particles. In general, we have $x_{i}=R+w_{1,i}\xi_{1}+w_{2,i}\xi_{2}$. Moreover, $\mathrm{d}^{4}x_{1}\,\mathrm{d}^{4}x_{2}\,\mathrm{d}^{4}x_{3}=|J^{-1}|^{4}\mathrm{d}^{4}R\,\mathrm{d}^{4}\xi_{1}\,\mathrm{d}^{4}\xi_{2}$. The conjugate four momenta of $R,\xi_{1},\xi_{2}$ are the total four momentum $P=p_{1}+p_{2}+p_{3}$ and a pair of Jacobi four momenta $q_{1},q_{2}$. We have $p_{1}\cdot x_{1}+p_{2}\cdot x_{2}+p_{3}\cdot x_{3}=P\cdot R+\xi_{1}\cdot q_{1}+\xi_{2}\cdot q_{2}$. In general, we find $p_{i}=c_{i}P+$ some linear combinations of $k_{1}$ and $k_{2}$, where $c_{1}+c_{2}+c_{3}=1$. Finally, $\mathrm{d}^{4}p_{1}\mathrm{d}^{4}p_{2}\mathrm{d}^{4}p_{3}=|J|^{4}\mathrm{d}^{4}P\,\mathrm{d}^{4}q_{1}\,\mathrm{d}^{4}q_{2}$.
[…]
@@
$\gamma\frac{\mathrm{d}N_{\mathrm{A}}}{\mathrm{d}^{3}P}=\frac{S_{\mathrm{A}}}{(2\pi)^{20}}|J^{-1}|^{4}\int\mathrm{d}^{4}P\int\mathrm{d}^{4}q_{1}\int\mathrm{d}^{4}q_{2}\int\mathrm{\penalty 10000\ d}^{4}R\int\mathrm{d}^{4}\xi_{1}\,\int\mathrm{d}^{4}\xi_{2}\int\mathrm{d}^{4}R^{\prime}\int\mathrm{d}^{4}\xi_{1}^{\prime}\int\mathrm{d}^{4}\xi_{2}^{\prime}\\ \times e^{-iP_{\mathrm{A}}\cdot(R-R^{\prime})}\varphi_{\mathrm{A}}^{*}\left(\xi_{1}^{\prime},\xi_{2}^{\prime}\right)\varphi_{\mathrm{A}}\left(\xi_{1},\xi_{2}\right)\,e^{iP\cdot\left(R-R^{\prime}\right)+iq_{1}\cdot\left(\xi_{1}-\xi_{1}^{\prime}\right)+iq_{2}\cdot\left(\xi_{2}-\xi_{2}^{\prime}\right)}\,\\ \times W_{\mathrm{A}}\left(p_{1},p_{2},p_{3},\frac{x_{1}+x_{1}^{\prime}}{2},\frac{x_{2}+x_{2}^{\prime}}{2},\frac{x_{3}+x_{3}^{\prime}}{2}\right)\,.$ (57)
@@
$\gamma\frac{\mathrm{d}N_{\mathrm{A}}}{\mathrm{\penalty 10000\ d}^{3}P}=\frac{S_{\mathrm{A}}}{(2\pi)^{16}}|J^{-1}|^{4}\int\mathrm{d}^{4}q_{1}\int\mathrm{d}^{4}q_{2}\int\mathrm{\penalty 10000\ d}^{4}S\int\mathrm{d}^{4}\xi_{1}\,\int\mathrm{d}^{4}\xi_{2}\int\mathrm{d}^{4}\xi_{1}^{\prime}\int\mathrm{d}^{4}\xi_{2}^{\prime}\\ \times\varphi_{\mathrm{A}}^{*}\left(\xi_{1}^{\prime},\xi_{2}^{\prime}\right)\varphi_{\mathrm{A}}\left(\xi_{1},\xi_{2}\right)\,e^{iq_{1}\cdot\left(\xi_{1}-\xi_{1}^{\prime}\right)+iq_{2}\cdot\left(\xi_{2}-\xi_{2}^{\prime}\right)}\,\\ \times W_{\mathrm{A}}\left(p_{1},p_{2},p_{3},\frac{x_{1}+x_{1}^{\prime}}{2},\frac{x_{2}+x_{2}^{\prime}}{2},\frac{x_{3}+x_{3}^{\prime}}{2}\right)\Bigr|_{P=P_{\mathrm{A}}}\,.$ (58)
For the moment everything is exact. We will now approximate this relation in order to obtain the three-dimensinal relation given in Eq. (4). In the following, we introduce the notation $S=(t_{S},\bm{S})$, $R=(t_{R},{\bm{R}})$, $\xi_{i}=(\tau_{i},\bm{\xi}_{i})$, $i=1,2$, and $x_{i}=(t_{i},{\bm{x}}_{i})$, $i=1,\ldots,3$, and similarly for the primed quantities. Clearly $t_{S}=(t_{R}+t_{R}^{\prime})/2$ and $t_{i}=t_{R}+$ linear combination of $\tau_{1}$ and $\tau_{2}$. For more details, see Ref. [23]. The approximations we consider are:
@@
    $\int{dq_{1}^{0}\over 2\pi}{dq_{2}^{0}\over 2\pi}e^{iq_{1}^{0}(\tau_{1}-\tau_{1}^{\prime})+iq_{2}^{0}(\tau_{2}-\tau_{2}^{\prime})}=\delta(\tau_{1}-\tau_{1}^{\prime})\delta(\tau_{2}-\tau_{2}^{\prime})\$(59)